### v1

# 1] QuantScape Inc. QuantScape Inc., 4-11-18, Manshon-Shimizudai, Meguro, Tokyo, 153-0064, Japan

Abstract
KAN: Kolmogorov–Arnold Networks is a novel multi-layer neuromorphic network that can exhibit higher accuracy than a neural network. It can learn and predict more accurately than neural networks with a smaller number of parameters, and many research groups worldwide have adopted it. As a result, many types of applications have been proposed. This network can be used as a generator solely or with a Generative Adversarial Network; however, KAN has a slower speed of learning than neural networks for the number of parameters. Hence, it has not been researched as a generator. Therefore, we propose a novel Generative Adversarial Network called Generative Adversarial Variational Quantum KAN that uses Variational Quantum KAN as a generator. This method enables efficient learning with significantly fewer parameters by leveraging the computational advantages of quantum circuits and their output distributions. We performed the training and generation task on MNIST and CIFAR-10, and revealed that our method can achieve higher accuracy than neural networks and Quantum Generative Adversarial Network with less data.
keywords:
Quantum machine learning , Born Machine , Quantum simulation , Noise
1 Introduction
Quantum computers are promised to solve some problems that classical computers are not able to solve since Lichard P Feynman proposed in 1982 [feynman_simulating_1982]. Some algorithms that can solve problems much faster than classical computers are reported by some groups, such as Grover algorithm [2003quant.ph…1079L], Quantum Fourier transform [PRXQuantum.4.040318] and its applications [2024arXiv241004435I, 2023arXiv230504908M], and Shor algorithms [2023arXiv230609122S].
However, quantum computers must avoid any types of noise to keep the fidelity of calculations rather than implementation [2024npjQI..10...18D] of required gates and circumvent noise is hard due to the lack of fault-tolerant architecture; hence, today‘s openly available quantum computers are mainly Noisy Intermediate-Scale Quantum(NISQ) devices.
In streamlining, quantum computing algorithms, particularly VQAs, have seen significant development in recent years. Foundational work by Aspuru-Guzik and colleagues [Kassal2011] led to algorithms such as VQE [McClean_2016], Adaptive VQE [Grimsley2019], and Multiscale Contracted VQE (MCVQE) [2019arXiv190608728P], so on [2021arXiv210501141W, 2021arXiv210902009W].
These algorithms are well-suited for NISQ devices and have been applied to quantum machine learning tasks[2014PhRvL.113m0503R, 2019QS&T....4a4001K, 2019Natur.567..209H, 2022PhRvA.106b2601A, 2022arXiv220211200K, 2021PhRvP..16d4057B, 2020PhRvL.125j0401W, 2022arXiv220608316Y, PhysRevA.98.032309] , demonstrating the potential of quantum approaches in enhancing AI models.
KAN: Kolmogorov–Arnold Networks is also one of neuro morphic multi-layer network which is emulated on the area of VQAs; Quantum KAN [2024arXiv241004435I] emulates KAN fully on fault-tolerant quantum computers by block encoding, Variational Quantum KAN [Wakaura_VQKAN_2024] emulates KAN by variational algorithms, and Kolmogorov-Arnold Network for Quantum Architecture Search [2024arXiv241004435I] takes advantage of KAN for assembling quantum circuit. They are more accurate than Quantum Neural Networks, hence, they are coming to be studied all over the world recently [2024arXiv241106727C, 2024arXiv241118165H], [2024arXiv241203710K, 2024arXiv241008452B, 2024arXiv241114902K, 2024arXiv241007446J, 2024arXiv240800273T]. However, KAN requires the number of parameters equivalent to the digital length of output data, the same as a Neural Network. Besides, computational time is parabolic for the number of parameters [Wakaura_VQKAN_2024]. A method to long data without increasing parameters is needed to apply quantum KAN families for complex tasks. Therefore, we established the method to emerge the long data by Variational Quantum KAN with support of the batch method and Generative Adversarial Network so-called GAVQKAN: Generative Adversarial Variational Quantum KAN. We performed the training of GAVQKAN and generated the test data on MNIST, CIFAR-10, and Fashion-MNIST datasets compared with Quantum GAN, classical Neural Network, and Quantum Reservoir GAN. As a result, we revealed that GAVQKAN can generate the MNIST, CIFAR-10 and Fashion-MNIST datasets accurately with less data than Quantum GAN and classical Neural Network. Section 1 is the introduction, Section 2 describes the method in detail of GAVQKAN, Section 3 describes the result of training and generating the pictures, Section 4 is the discussion of our results, and Section 5 is the concluding remark.
2 Method
In this section, we describe the details of GAVQKAN.
GAVQKAN is the Quantum GAN that uses KAN on ansatz and generates the data obeying the manner of Born Machine. GAN is one of the generative models that combines two networks or models in machine learning. The discriminator distinguishes genuine data from fake data generated by the generator network, and the generator network makes the fake data pass the discriminator’s judgment as genuine data. Both are optimized to accomplish the assigned task, and they are expressed as an equation,
$\min_{G}\max_{D}V_{\text{\tiny GAN}}(D,G)=\mathbb{E}_{\bm{x}\sim p_{\text{data}}(\bm{x})}[\log D(\bm{x})]+\mathbb{E}_{\bm{z}\sim p_{\bm{z}}(\bm{z})}[\log(1-D(G(\bm{z})))]$ (1)
Then, $D(\bm{x})$ is the probability that the data $\bm{x}$ is distinguished correctly for the boolean label of the data $p_{\text{data}}(\bm{x})$ on discriminator $D$, and $G(\bm{z})$ is the data made by generator $G$ from latent vector $\bm{z}$, respectively. The boolean label of the data $\sim p_{\text{data}}(\bm{x})$ becomes 0 if the data is fake and 1 if genuine, and $p_{\bm{z}}(\bm{z})$ is 1 for all cases. In simple form, it is expressed as,
$V_{\text{\tiny GAN}}(D,G)$ $=$ $L_{D}+L_{G}$ (2)
$L_{D}$ $=$ $p\log D(\bm{x})+(1-p)\log D(G(\bm{z}))$ (3)
$L_{G}$ $=$ $-p\log(1-D(G(\bm{z})))$ (4)
which $p$ is the boolean label. They are loss functions to be minimized, hence, the generator loss function is referred as
$l_{G}=-\log(D(G(\bm{z}))),$ (5)
to be maximized.
We generate the $G(\bm{z})$ by the Variational Quantum KAN using the latent vector $z$ as noise. VQKAN is the variational quantum algorithm version of KAN, a multi-layer network based on the connection of synapses in neurons. First, initial state $\mid\Psi_{ini}(\bm{z})\rangle$ is $\prod_{j=0}^{N_{q}-1}Ry^{j}(\bm{z})\mid 0\rangle^{\otimes N_{q}}$ for each input $m$. $Ry^{j}(\theta)$ is $\theta$ degrees angle rotation gate for y-axis on qubit $j$. $_{n}{\bf x}$ is the input vector at layer n which $_{1}{\bf x}=\bm{z}$. For VQKAN, $\phi_{jd}^{n}(_{n}{\bf x})$ is the gate of the angle.
$\phi_{jd}^{n}(_{n}{\bf x})=\sum_{i\in\{0,dim(_{n}{\bf x})\}}^{0}2acos(E_{f}(_{n}x_{i})+\sum_{s=0}^{N_{g}-1}\sum_{l=0}^{N_{s}-1}c_{s}^{njd}B_{l}(_{n}x_{i})).$ (6)
, which $n$ is the index of the layer, $N_{g}$ is the number of grids for each gate, $N_{s}$ is the number of splines, respectively. Then, $c_{s}^{njd}$ and $B_{l}(_{n}x_{i})$ are the parameters to be trained, initialized into 0 and B spline functions at layer n whose domains are $[0,1]$, respectively, the same as classical KAN. $_{n}{\bf x}$ is the input vector at layer n, $d$ is the index of the depth in the single layer, and $j$ is the index of qubits, respectively. $E_{f}(_{n}x_{i})=_{n}x_{i}/(exp(-_{n}x_{i})+1)$ is the Fermi-Dirac expectation energy-like value of the distribution. The component of $_{n}{\bf x}$ is the expectation value of the given observable for the calculated states of qubits. The layers are the combined of para metric gates called ansatz. In detail the single layer $\Phi_{n}^{G}=\prod_{d=0}^{N_{d}-1}(\prod_{j=0}^{N_{q}-1}Ry^{j}(\phi_{jd}^{n}(_{n}{\bf x}_{j}))\prod_{j=0}^{N_{q}-2}CZ_{j,j+1})$ as shown in Fig.1 with the abstract picture of the flow of GAVQKAN. The entire ansatz is as follows.
$\mid\Psi(_{1}{\bf x})\rangle=\prod_{n=1}^{num.~of~layers~N_{l}}{\Phi}_{n}^{G}M\mid\Psi_{ini}(_{1}{\bf x})\rangle$ (7)
Then, $M$ indicates the measurement of all qubits and deriving distribution of probability as a $2^{N_{q}}$ - length vector at a maximum length to make $_{n+1}{\bf x}$. $_{n}{\bf x}_{j}$ for $n>1$ is the average of the $N_{q}$ equally separated segments of output data.
Figure 1: The simplified illustration of our method. Each qubit is initialized by a latent vector $z$. Qubits are calculated by Ry gates, and blended by control - Z gate, respectively by layer. Quantum circuits emerge from probability distributions as patches, and they are unified into one data. Loss functions are two cross entropy calculated from genuine and generated data.
The condition of convergence is the default of scipy for all methods. We assume $N_{l}=1,N_{q}=8,N_{g}=8$ and $N_{s}=4(tr+2)$ for the number of trials $tr$, respectively. $G(\bm{z})$ is the combined probability distribution from 4 patches. The length of probability distribution we use as the patch is the quarter of the length of the entire image as a vector. We use stochastic gradient descent as an optimizer on both the discriminator and the generator, and the learning rates are 0.1 and 0.001, respectively. We train once on an iteration on one image. The number of samples to process on the GAVQKAN of an iteration is fixed to 1. All calculations are performed using PyTorch and Google Colab. We use PennyLane as a quantum simulator.
3 Result of numerical simulations
In this section, we describe the results of generating the pictures on the MNIST and CIFAR10 dataset. First, we compared the results on GAVQKAN, Quantum GAN (QGAN), and Classical Neural Network (CNN) using loss functions of cross-entropy on 16 $\times$ 16-sized handwritten single numbers for 1000 samples, respectively. QGAN [2021PhRvP..16b4051H] is simulated by PennyLane. Number of qubit is 8 and the depth is 6, and ansatz is $\prod_{j=0}^{N_{q}-1}Ry(z_{j})\prod_{k=0}^{N_{d}-1}\prod_{j=0}^{N_{q}-1}Ry(\theta_{j,k})\prod_{j=0}^{N_{q}-2}CZ_{j,j+1}$ for noise and parameters for each qubit. CNN is performed by a 6-layer network consisting of layers sized [ 256, 1024, 2048, 6272, 6272, 256 ]. Discriminator is CNN consists of 3 layers sized [ $N_{data}$, $2\sqrt{N_{data}}$, 1 ] for the data length $N_{data}$ with the rectified linear unit function between marginal layers and Sigmoid at the end.
We show the average of the loss function for 10 attempts of the discriminator and generator of GAVQKAN, QGAN, and CNN for the number of epochs using cross-entropy in Fig.2, and the generated picture of 0 in Fig.3, respectively. The loss functions of generator GAVQKAN became over 6 the firstest, and QGAN is the second firstest. In contrast, the loss function of the discriminator of GAVQKAN became 0 the slowest in all, and QGAN is the second slowest. The generated pictures of GAVQKAN are close to genuine data, even though the number of iterations is under 100. Although, generated pictures of QGAN became closer to genuine data than those of GAVQKAN when the number of iterations is over 400. Sliced Wasserstein Distance (SWD) also exhibits the accuracy and their tendency.
Figure 2: The number of iterations v.s. the average of the loss function for 10 attempts of the discriminator $L_{D}$ and generator $l_{G}$ of GAVQKAN, QGAN, and CNN using cross-entropy as a loss function. Error bars indicate the standard deviation from the average, and each bar is sampled every 10 points, moving 2 points from the previous data.
Figure 3: The generated picture of handwritten numbers by GAVQKAN, QGAN, and CNN using cross-entropy as a loss function for the number of iterations. The picture by GAVQKAN is generated by fixed noise and input.
Figure 4: The number of iterations v.s. the average of the MSE and SWD for 5 attempts of GAVQKAN, QGAN, and CNN using cross-entropy as a loss function on the training and generating the MNIST dataset.
We also show the MSE and SWD of three methods of the first 8 pictures for fixed noise in Fig.4. SWDs are calculated by Python Optimal Transport [JMLR:v22:20-451] on the first 8 pictures in the index of the original dataset. SWD of GAVQKAN in the range of the number of iterations below 400 is the smallest in all, and the second smallest over 400. GAVQKAN can generate the data accurately by the smallest numbers of data in GAVQKAN, QGAN, and CNN. However, the accuracy itself is not high compared to QGAN. In addition, MSE of GAVQKAN is the second smallest in average in all. The loss function of the discriminator of GAVQKAN has the smallest standard deviation in all; hence, the learning ratio is not supposed to be the optimum value. Besides, the time for calculations is 6.5569 times larger than QGAN and 11.4687 times larger than CNN, respectively. The number of parameters and ansatz must be surveyed further to improve both calculation time and accuracy. Next, we compared the result of generating a 22 $\times$ 22 size picture on the CIFAR10 dataset. We used only 1000 data from index 0. Both pictures are compressed from 32 $\times$ 32 size original pictures into 22 $\times$ 22 size and converted to gray scale. The number of qubits on QGAN is 8, and the size of the network on CNN is 8 times that of training on 0. We show the average of the loss function for 10 attempts of the discriminator and generator of GAVQKAN, QGAN, and CNN for the number of epochs using cross-entropy in Fig.5, and the generated picture of the first 8 pictures for fixed noise in Fig.6, respectively. The average of GAVQKAN did not grow compared to QGAN and other cases of training, even though generated pictures have the smallest SWD in all, also shown in Fig.7. We show the MSE and SWD of three methods in Fig.7. The loss functions of the generator of GAVQKAN became the largest in all, and that of QGAN is the second largest, the same as MNIST. In contrast, the loss functions of generators of GAVQKAN became the smallest and suppressed the slowest in all. SWD of GAVQKAN in the range of the number of iterations below 400 is the smallest in all, and the second smallest over 400, the same as MNIST. Besides, the time for calculations is 6.4084 times larger than QGAN, 19.6673 times larger than CNN, and 0.7437 times larger than GAVQKAN, respectively. GAVQKAN takes less time than QRGAN because GAVQKAN emerges with longer data than MNIST. GAVQKAN exhibits high accuracy with a small amount of data, as in the case of MNIST.
Figure 5: The number of iterations v.s. the average of the loss function for 5 attempts of the discriminator and generator of GAVQKAN, QGAN, and CNN using cross-entropy as a loss function on the training and generating the CIFAR10 dataset.
Figure 6: The generated picture of the first 8 data by GAVQKAN, QGAN, and CNN using the cross-entropy as a loss function. The pictures in the upper right are the original pictures from 0 to 9 on the index.
Figure 7: The number of iterations v.s. the average of the MSE and SWD for 5 attempts of GAVQKAN, QGAN, and CNN using cross-entropy as a loss function on the training and generating the CIFAR10 dataset.
4 Discussion
In this section, we discuss the reproduction property and the accuracy for the noisy input of GAVQKAN. First, we discuss the reproduction property of the output of GAVQKAN for data on Fashion-MNIST. We generated the first 8 data learning 1000 data using GAVQKAN, and calculated the loss functions of discriminator, the loss functions of generator, MSE and SWD for seed coefficient of random numbers 42,2,10,13,0,3407,7120,10000,11111,16384,17171,130000,14480,11668,50000142,2,10,13,0,3407,7120,10000,11111,16384,17171,130000,14480,11668,500001 and $620000$ and calculated t and p values by Bonferroni test. Data are all original size ( $16\times 16$ ). We show the average of the loss function, MSE, and SWD for 16 seeds of GAVQKAN for the number of epochs using cross-entropy in Fig.8, and the generated picture of the first 8 pictures for fixed noise in Fig.9, respectively. The generator learned and was enlarged by the progress of learning. The average of SWD is nearly 0.35, the same as the result of CIFAR-10. According to the result of the Bonferroni test on SWD for all iterations, at least 9 seeds are clearly different from the given seed ( shown in Table.1 ). The significance is $0.05/240$, hence, the two seeds that have p value over 0 differ negligibly. SWD for different seeds has a little difference compared to QRGAN.
Figure 8: The number of iterations v.s. the average of the loss function, MSE, and SWD for 16 seeds of the discriminator and generator of GAVQKAN using cross-entropy as a loss function.
Figure 9: The generated picture of the first 8 data for seed 0 by GAVQKAN using the cross-entropy as a loss function.
Table 1: The value of t and p values between two seeds on Fashion-MNIST of SWD of GAVQKAN using cross-entropy as a loss function.
t value 42 2 10 13 0 3407 7120 10000 11111 16384 17171 130000 14480 11668 500001 620000
42 0 0.1605 4.2064 -0.6913 4.551 -0.0434 2.1098 0.2537 -1.2875 1.1109 0.789 1.7223 -0.631 4.2608 4.4995 -0.0683
2 -0.1605 0 3.9293 -0.8319 4.26 -0.2006 1.8909 0.0854 -1.4117 0.9199 0.6094 1.5178 -0.7715 3.9721 4.2177 -0.2251
10 -4.2064 -3.9293 0 -4.8694 0.2988 -4.1911 -2.0958 -3.974 -5.4579 -3.09 -3.3596 -2.4206 -4.7491 -0.0576 0.3315 -4.2253
13 0.6913 0.8319 4.8694 0 5.2185 0.6387 2.7868 0.9467 -0.5935 1.7939 1.4659 2.3921 0.0501 4.94 5.1549 0.6155
0 -4.551 -4.26 -0.2988 -5.2185 0 -4.5299 -2.4174 -4.3181 -5.8125 -3.422 -3.6898 -2.7403 -5.091 -0.3653 0.039 -4.5654
3407 0.0434 0.2006 4.1911 -0.6387 4.5299 0 2.1237 0.2937 -1.2268 1.1388 0.8211 1.7417 -0.58 4.2427 4.4807 -0.0245
7120 -2.1098 -1.8909 2.0958 -2.7868 2.4174 -2.1237 0 -1.8668 -3.379 -0.9963 -1.2919 -0.356 -2.6964 2.0941 2.4081 -2.1533
10000 -0.2537 -0.0854 3.974 -0.9467 4.3181 -0.2937 1.8668 0 -1.5454 0.8631 0.5427 1.4809 -0.8824 4.0231 4.2705 -0.3193
11111 1.2875 1.4117 5.4579 0.5935 5.8125 1.2268 3.379 1.5454 0 2.3873 2.0526 2.9765 0.6352 5.5435 5.7376 1.2049
16384 -1.1109 -0.9199 3.09 -1.7939 3.422 -1.1388 0.9963 -0.8631 -2.3873 0 -0.3077 0.6259 -1.7179 3.1146 3.3929 -1.1661
17171 -0.789 -0.6094 3.3596 -1.4659 3.6898 -0.8211 1.2919 -0.5427 -2.0526 0.3077 0 0.9219 -1.3956 3.3898 3.6567 -0.8474
130000 -1.7223 -1.5178 2.4206 -2.3921 2.7403 -1.7417 0.356 -1.4809 -2.9765 -0.6259 -0.9219 0 -2.3091 2.4264 2.7256 -1.77
14480 0.631 0.7715 4.7491 -0.0501 5.091 0.58 2.6964 0.8824 -0.6352 1.7179 1.3956 2.3091 0 4.8137 5.0322 0.557
11668 -4.2608 -3.9721 0.0576 -4.94 0.3653 -4.2427 -2.0941 -4.0231 -5.5435 -3.1146 -3.3898 -2.4264 -4.8137 0 0.3973 -4.2782
500001 -4.4995 -4.2177 -0.3315 -5.1549 -0.039 -4.4807 -2.4081 -4.2705 -5.7376 -3.3929 -3.6567 -2.7256 -5.0322 -0.3973 0 -4.5152
620000 0.0683 0.2251 4.2253 -0.6155 4.5654 0.0245 2.1533 0.3193 -1.2049 1.1661 0.8474 1.77 -0.557 4.2782 4.5152 0
p value 42 2 10 13 0 3407 7120 10000 11111 16384 17171 130000 14480 11668 500001 620000
42 1 0.8725 0 0.4894 0 0.9654 0.035 0.7998 0.1981 0.2667 0.4302 0.0852 0.5281 0 0 0.9456
2 0.8725 1 0.0001 0.4056 0 0.841 0.0588 0.932 0.1582 0.3577 0.5423 0.1292 0.4405 0.0001 0 0.8219
10 0 0.0001 1 0 0.7651 0 0.0362 0.0001 0 0.002 0.0008 0.0156 0 0.9541 0.7403 0
13 0.4894 0.4056 0 1 0 0.5231 0.0054 0.3439 0.5529 0.073 0.1428 0.0168 0.96 0 0 0.5383
0 0 0 0.7651 0 1 0 0.0157 0 0 0.0006 0.0002 0.0062 0 0.715 0.9689 0
3407 0.9654 0.841 0 0.5231 0 1 0.0338 0.769 0.22 0.2549 0.4117 0.0817 0.562 0 0 0.9805
7120 0.035 0.0588 0.0362 0.0054 0.0157 0.0338 1 0.0621 0.0007 0.3192 0.1965 0.7219 0.0071 0.0364 0.0161 0.0314
10000 0.7998 0.932 0.0001 0.3439 0 0.769 0.0621 1 0.1224 0.3882 0.5874 0.1388 0.3777 0.0001 0 0.7495
11111 0.1981 0.1582 0 0.5529 0 0.22 0.0007 0.1224 1 0.0171 0.0402 0.003 0.5254 0 0 0.2284
16384 0.2667 0.3577 0.002 0.073 0.0006 0.2549 0.3192 0.3882 0.0171 1 0.7584 0.5314 0.086 0.0019 0.0007 0.2437
17171 0.4302 0.5423 0.0008 0.1428 0.0002 0.4117 0.1965 0.5874 0.0402 0.7584 1 0.3567 0.163 0.0007 0.0003 0.3969
130000 0.0852 0.1292 0.0156 0.0168 0.0062 0.0817 0.7219 0.1388 0.003 0.5314 0.3567 1 0.021 0.0153 0.0065 0.0769
14480 0.5281 0.4405 0 0.96 0 0.562 0.0071 0.3777 0.5254 0.086 0.163 0.021 1 0 0 0.5776
11668 0 0.0001 0.9541 0 0.715 0 0.0364 0.0001 0 0.0019 0.0007 0.0153 0 1 0.6912 0
500001 0 0 0.7403 0 0.9689 0 0.0161 0 0 0.0007 0.0003 0.0065 0 0.6912 1 0
620000 0.9456 0.8219 0 0.5383 0 0.9805 0.0314 0.7495 0.2284 0.2437 0.3969 0.0769 0.5776 0 0 1
Next, we surveyed the accuracy of GAVQKAN for the case a : depth: 6, layer: 1, b : canonical ansatz, depth, layer: 1 , c : radial base function, d : radial + canonical ansatz, e : radial + depth: 8, layer: 1, f : radial + num. qubits: 10, depth: 6, layer: 1, g : radial + depth: 6, layer: 2, h : radial + Adam, i : QGAN, depth: 8, j : KAN [ 256,256,512,1024,256 ] , and discussed the method to improve it on the MNIST dataset. The canonical ansatz is the ansatz of paper, radial base function indicates the case uses the radial base function instead of B spline function on eq.6, and Adam indicates the optimizer is the Adam method. The canonical ansatz saves time for calculation and improves the accuracy a little compared to our ansatz, as shown in Fig. 10 and Table. 2. The radial base function instead of B spline function on KAN layers saved prominent time for calculation, prominently in exchange for only about 0.02 SWD on average, Fig. 10 and Table. 2. Cases e and f improved the accuracy and reached the accuracy of QGAN including case i. Adam optimizer saves time for calculation further. However, the loss functions converged far faster than all other cases. Case a-h are all lowered a little or nearly never changed differently from cases i and j. We suppose that the learning rates of the discriminator and generator are inadequate to train them enough. Tuning them may be read to improve accuracy. Besides, applying APA(Adaptive Pseudo Augmentation) [2021arXiv211106849J] and other improved GAN [9770060, 2021arXiv211101118K] can improve the accuracy. KAN itself requires a large number of iterations for convergence. The robustness of VQKAN against overfitting may lower the speed of convergence.
Figure 10: The number of iterations v.s. the average of the SWD of 10 cases on the MNIST and CIFAR10 dataset using cross-entropy as a loss function.
Table 2: The speed of calculations on each case for 1000 iterations. The time for calculation of case a is $t_{0}$.
case speed ( $t_{0}/t_{n}$ )
a 1
b 1.3134
c 3.6527
d 3.6966
e 2.6299
f 3.1761
g 1.8904
h 9.4137
i 4.7942
j 51.749
5 Concluding remarks
In this paper, we revealed that GAVQKAN has more accuracy than QGAN, CNN, and classical KAN on the MNIST and CIFAR10 datasets for small amounts of data. This result may contribute to the generation of large data by a larger number of qubits. Besides, the accuracy for the number of parameters is better than classical KAN. However, the speed of convergence is slower than that of CNN and classical KAN. Improving it is the next problem. Our work is only on simulators; hence, calculating on real quantum devices is also the next plan.